### v1

@@ IV.2 Emulation with fixed $H$ and sources
The primary purpose is to understand the spectrum of a fixed $H$, as compressed in the emulator subspace, and its dependence on the $N_{b}$ and the locations of $E^{tr}_{\alpha}$ in the complex $E$ plane. The latter dependence suggests a general choice of $E^{tr}_{\alpha}$ suitable for continuum physics. Another main observation is the similarities between the CEE and the (near)-optimal rational approximation.
We choose $|{S}\rangle=|\tilde{{S}}\rangle=V|P_{rel},\,\ell\rangle$ with $|P_{rel},\,\ell\rangle$ as the incoming plane wave ($P_{rel}=\sqrt{M_{N}E_{rel}}$), so that $\mathcal{A}$ corresponds to the non-Born term of $T$-matrix (labeled as $T_{nonBorn}$). The sources and $H$ are fixed by fixing $E_{rel}$ $\lambda$ and $\Lambda_{2}$. Here, $\Lambda_{2}=200$ MeV, and $\lambda$ is tuned to get a desired bound or resonance state in the $H$’s spectrum. About the training points, $\mathrm{Im}\,E^{tr}_{\alpha}$ is fixed to one or two values, while $\mathrm{Re}\,E^{tr}_{\alpha}$ are sampled on an even mesh in a $[-20,20]$ MeV interval (see the black lines in e.g., Fig. 3).
Figure 3: The relative error for emulated $\mathcal{A}$ in the complex $E$ plane. In each sub-figure, from left to right, $N_{b}$ increases from 3 to 23, while from top to bottom $\mathrm{Im}\,E^{tr}_{\alpha}$ decreases from 30 to 3 MeV. As discussed in the main text, $E^{tr}_{\alpha}$ are evenly separated along the black lines in each panel. The two sub-figures differ in the arrangement of the training points.
First, we aim to understand the CEE via its errors, which is also the primary approach for studying the (near)-optimal rational approximation Trefethen (2023) [^16] For example, the pole distribution pattern in the (near)-optimal rational approximation can be understood as the result of minimizing the approximation errors Trefethen (2023).. Fig. 3 shows such relative errors for the emulated $s$-wave $\mathcal{A}(E)$s in the complex $E$ plane. Here, the two-body binding energy $B_{2}$ is $10$ MeV and $E_{rel}=10$ MeV in the sources. In the left sub-figure, the training points are lined on one black line above the real axis, but in the right, they are distributed on two black lines corresponding to a complex conjugate pair. The latter setup was mentioned previously at the end of Sec. III.2. Four emulations are contrasted within each sub-figure, based on different combinations of $\mathrm{Im}\,E^{tr}_{\alpha}$ and $N_{b}$ values.
@@
The general theme of the left sub-figure is that the errors are in the same order as those of the training calculations in the interpolation region (colored dark blue) between the training points or close to them. The errors increase when evaluating $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ (i.e., the emulated $\mathcal{A}$) further away. The orange dots representing the largest emulation errors are the locations of the DC poles of $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ because $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ diverges at those locations but the exact results are finite. Those orange dots visually trace out a BCT of the exact $\mathcal{A}$ that is significantly below the real $E$ axis, even though the high-fidelity calculation with the Hermitian $H$ operator would have a BCT along the real axis (see, e.g., Fig. 1). I.e., those DC poles form a proxy for the BCT of the exact $\mathcal{A}$ that the emulator approximates. This is the same as how the (near)-optimal rational approximation works. We now consider $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ effectively having the discretized BCT.
As it turns out, choosing an exact $T$-matrix to be compared with the emulated one could be subtle in the 4th quadrant of the complex $E$ plane. Note in the current discussion, we define the Riemann sheet of the exact $\mathcal{A}$ according to the BCT defined on the real axis (e.g., as in the analytic formulas in Appendix B). The exact result in the quadrant 1 to 3 is the one on the 1st sheet. In quadrant 4, above the new discretized BCT, the exact result used for computing emulation errors is the $\mathcal{A}$ defined on the 2nd sheet; however, below the new BCT, the exact result is the one on the 1st sheet. The ambiguity could arise for the Riemann sheet assignment in the region close to the BCT poles. Here, we connect the adjacent points by a straight line to form a continuous BCT curve, which determines the Riemann sheet of the exact $\mathcal{A}$. The well-behaved error plot near the new BCT shows the simple procedure is sufficient in the current cases; otherwise, sharp edges could show up in the region away from the discretized BCT.
@@
In the right sub-figure, where the emulators are trained on both $E^{tr}_{\alpha}$ and their complex conjugates, the dark orange dots, i.e., the discretized BCT, are back on the real axis in all the panels. This reinforces the relation between the error pattern and the BCT location. The error pattern shares the same mirror symmetry with respect to the real axis as the locations of the training points. Another possibility for the discretized BCT, while respecting the same symmetry, is to have poles distributed symmetrically on both sides of the real axis. As argued in the Appendix A, this case would have more significant emulation errors in general than the existing one with a single discretized BCT on the real axis.
@@
The difficulty in describing continuum physics with BCT on the real axis is also exposed clearly in these panels. Reducing $\mathrm{Im}\,E^{tr}_{\alpha}$ and increasing $N_{B}$ doesn’t improve $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ on the real axis, although it does improve in the region of the bound states. Therefore, we should follow the training point setup used in the right sub-figure to infer the discrete excited states from CEE; however, in the following parts of the paper focusing on continuum physics, we distribute the training points as in the left sub-figure.
@@
Figure 4: The left panels show the compressed spectra. The right ones plot the residues of the poles (see Eq. (36)) vs. the absolute values of the corresponding eigenvalues. Their insets show the absolute values of the eigenenergies in the $\log$ scale. The physical (bound) state is marked with a black “$+$” in the left panels and the insets of the right panels. From the top and bottom rows, $N_{b}$ increases from $3$ to $23$. Each row compares two emulation results with different $\mathrm{Im}\,E^{tr}_{\alpha}$.
@@
Figure 4 shows the information about the compressed spectra and the residues of the corresponding poles in $\begin{bmatrix}\mathcal{A}(E)\end{bmatrix}$s (see Eq. (36)) from the emulators shown in the left sub-figure of Fig. 3. From top to bottom, $N_{b}$ increases from 3 to 23. In each panel, the results from two emulators with different $\mathrm{Im}\,(E_{\alpha}^{tr})$ are compared.
@@
The left panels again show the BCT pattern of the DC state eigenenergies, while the bound state (not visible in Fig. 3) is separated from the DC states. This feature is similar to the eigenenergy distributions in the complex scaling calculations Myo et al. (2014), but the number of eigenvalues here is much smaller than that in a typical complex scaling calculation. Moreover, in the complex scaling method, the angle between the rotated BCTs and the real axis is a control parameter, the maximum of which could be limited due to numeric difficulties. In the CEE, the moved BCTs, which emerge from spectrum compression, could be affected by the $\mathrm{Im}\,E_{\alpha}^{tr}$: the smaller the $\mathrm{Im}\,E_{\alpha}^{tr}$ gets, the further away the BCTs are pushed from the real axis. Such behavior is reflected in the left panels.
@@
This could be understood intuitively: reducing $\mathrm{Im}\,E_{\alpha}^{tr}$ decreases the emulation errors below the real axis, and thus effectively, the failure region, where the discretized BCTs are located, is pushed further away from the real axis. Note the physical states, such as the bound state in the left panels, have less sensitivity. Therefore, this difference could be exploited to separate the physical states from the DC states, in addition to relying on the distribution pattern of the eigenenergies or poles. These two ways of separating states all trace back to the fact that the two types of states or poles represent the smooth and peak components separately in the $-\mathrm{Im}\,\mathcal{A}$ along the real axis. The physical states are independent of the BCT choice because $-\mathrm{Im}\,\mathcal{A}$ on the real axis and its decomposition into smooth and peak components are also independent. In contrast, the DC states can be redistributed when we redefine the BCT, as long as the physical and DC states are well separated. This point is further elaborated in Sec. V.
@@
The right panels in Fig. 4 plot the sizes of the pole residues ($W_{\beta}$ in Eq. (36)) against the absolute values of the eigenvalues. We see that incredibly close to the branch point (threshold), there could be poles with tiny residues when $N_{b}$ becomes too large, such as in the $N_{b}=23$ case. They could be the so-called Froissart doublets (or spurious poles, see Sec. 8.1 in Ref. Trefethen (2023)), whose contributions are negligible unless $E$ gets exceptionally close to the threshold (much smaller than $10^{-10}$ MeV here). If we know the thresholds from subsystem mass computations, we can separate and eliminate these spurious poles in Eq. (36). Or, we could apply the so-called greedy algorithm Sarkar and Lee (2021b); Quarteroni et al. (2016) to efficiently sample the training points in the parameter space and stop at a $N_{b}$ that is enough for spanning the solution subspace. For now, $N_{b}=3$ is under-spanning the subspace, while $N_{b}=23$ is over. Therefore, the effective dimension of the subspaces for both emulators is between $3$ and $23$.
@@
Each inset plots the absolute values of the eigenenergies on a logarithmic scale, with the bound state marked in black $+$. The distribution of the DC states is quasi-even on the log scale, with a gradual tapering towards both the threshold and the infinity, which is already visible in the left panels with linear scales. This distribution is similar to the tapered exponential clustering of poles in the (near)-optimal rational approximations discussed in Sec. II.3. We thus identify the first similarity suggesting a close connection between CEE’s dimension reduction efficiency and the optimality of the rational approximation.
@@
Also note that when reducing $\mathrm{Im}\,E_{\alpha}^{tr}$, the smallest distance between the BCT poles (or DC state eigenvalues)—excluding the spurious poles—and the threshold decreases; the distribution of poles or the eigenvalues also gets denser. The scale of that smallest separation controls to what proximity towards the threshold the emulator can be trusted because those close poles are responsible for the near-threshold behavior of $\begin{bmatrix}\mathcal{A}\end{bmatrix}$. The emulations and rational approximations eventually fail if $E$ gets too close to the threshold, as mentioned at the end of Sec. II.3.
@@
Of course, the emulation performance is also related to the number of effective poles (excluding the spurious poles), defined as $N_{\mathrm{eff}}$. As we can imagine, when $N_{b}$ increases beyond necessary, i.e., $N_{\mathrm{eff}}$ approaches the dimension of the solution subspace, the quality of near-threshold approximation saturates, and the smallest distance between the pole and branch point becomes “stuck”. This saturation behavior is further illustrated later in the Fig. 5. Therefore, by counting the number of symbols in the insets, i.e., the effective poles, for the $N_{b}=23$ cases, we see that the $\mathrm{Im}\,E^{tr}_{\alpha}=3$ subspace is higher dimensional than that of $\mathrm{Im}\,E^{tr}_{\alpha}=30$. This indicates $\mathrm{Im}\,E^{tr}_{\alpha}$ could be viewed as a continuum resolution indicator for the solution manifold mapped out by varying the $\mathrm{Re}\,E$ parameter. Such resolution, an intrinsic property of the solution manifold, can only be “measured” by the emulator when its $N_{\mathrm{eff}}$ saturates the subspace dimension.
@@
Figure 5: The emulation errors vs $\sqrt{N_{\mathrm{eff}}}$. The left panels show the emulation errors for the bound state pole location and its residues. The right ones show the absolute errors for $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ (1) on the real energy axis with different $E$ values marked in the legends, and (2) averaged over the interpolation region around the training energies (marked as “int.”). From the top to bottom row, $\mathrm{Im}\,E^{tr}_{\alpha}$ reduces from $30$ to $3$ MeV.
@@
Fig. 5 shows detailed information about the emulation errors vs $\sqrt{N_{\mathrm{eff}}}$ for the pole properties, including its locations and residues [^17] These residues could be related to the asymptotic normalization coefficients of the bound states Zhang et al. (2018). in the left panels. The right panels show the same information for $\begin{bmatrix}\mathcal{A}(E)\end{bmatrix}$ in the interpolation region (marked as “int” in the legend) and along the real energy axis with locations marked in the legend (considered as extrapolations). $\mathrm{Im}\,E_{\alpha}^{tr}$ decreases from the top to bottom panels. To get the data points for these plots, we run emulations with different $N_{b}$ and compute the corresponding $N_{\mathrm{eff}}$ as already described.
@@
We can see the $e^{-c\sqrt{N_{\mathrm{eff}}}}$ convergence behavior for all the quantities of our interests. Towards the largest $N_{\mathrm{eff}}$, our SVD-based regularization could also contribute to the emulation errors. Moreover, the extrapolation errors are generally more significant than the interpolation errors but still satisfy the same scaling law. It is also clear that reducing $\mathrm{Im}\,E_{\alpha}$ increases the maximum $N_{\mathrm{eff}}$ value we can achieve, which is essential for lowering both interpolation and extrapolation errors. This supports the reasoning about $N_{\mathrm{eff}}$ saturating the subspace dimension during the discussion of Fig. 4. The convergence behavior is the second similarity that suggests the connection between the CEE and (near)-optimal rational approximations.
@@
Now, we look into the $p$-wave channel to investigate emulation performance with the presence of a resonance. Here, $\Lambda_{2}=200$ MeV and $\lambda=0.8$ with a resonance located at $E=3.9-2.6\,i$ MeV. Figs. 6, 7, and 8 show the results, in parallel to the plotted results in the $s$-wave case. The key difference from the $s$-wave is that the physical pole is now located below the real axis and thus has bigger errors than in the previous $s$-wave results. However, all the other observations are the same as in the $s$-wave case.
@@
Figure 6: Similar to Fig. 3, but for the $p$-wave case with $\Lambda_{2}=200$ MeV and $\lambda=0.8$. The resonance pole is located at $3.9-2.6\,i$ MeV.
@@
Figure 7: Similar to Fig. 4, but for the $p$-wave case with $\Lambda_{2}=200$ MeV and $\lambda=0.8$. The resonance pole is located at $3.9-2.6\,i$ MeV.
Figure 8: Similar to Fig. 5 but for $p$-wave case with $\Lambda_{2}=200$ MeV and $\lambda=0.8$. The resonance pole is located at $3.9-2.6\,i$ MeV.
IV.3 Emulation in ($\mathrm{Re}\,E$, $\mathrm{Im}\,E$, $E_{rel}$, $\lambda$, $\Lambda_{2}$)

### v2

@@ IV.2 Emulation with fixed $H$ and sources
The primary purpose is to understand the spectrum of a fixed $H$, as compressed in the emulator subspace, and its dependence on the $N_{b}$ and the locations of $E^{\mathrm{tr}}_{\alpha}$ in the complex $E$ plane. The latter dependence suggests a general choice of $E^{\mathrm{tr}}_{\alpha}$ suitable for continuum physics. Another main observation is the similarities between the CEE and the (near-)optimal rational approximation.
I again choose $|{S}\rangle=|\tilde{{S}}\rangle=V|P_{\mathrm{rel}},\,\ell\rangle$ with $|P_{\mathrm{rel}},\,\ell\rangle$ as the incoming plane wave ($P_{\mathrm{rel}}=\sqrt{M_{N}E_{\mathrm{rel}}}$), so that $\mathcal{A}$ corresponds to the non-Born term of $T$-matrix (labeled as $T_{\mathrm{nonBorn}}$). The sources and $H$ are fixed by fixing $E_{\mathrm{rel}}$ $\lambda$ and $\Lambda_{2}$. Here, $\Lambda_{2}=200$ MeV, and $\lambda$ is tuned to get a desired bound or resonance state in the $H$’s spectrum. About the training points, $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ is fixed to one or two values, while $\mathrm{Re}\,E^{\mathrm{tr}}_{\alpha}$ are sampled on an even mesh in a [20,20][-20,20] MeV interval (see the black lines in e.g., Fig. 3).
Figure 3: The relative error for emulated $\mathcal{A}$ in the complex $E$ plane. In each sub-figure, from left to right, $N_{b}$ increases from 3 to 23, while from top to bottom $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ decreases from 30 to 3 MeV. As discussed in the main text, $E^{\mathrm{tr}}_{\alpha}$ are evenly separated along the black lines in each panel. The two sub-figures differ in the arrangement of the training points.
First, I aim to understand the CEE via its errors, which is also the primary approach for studying the (near-)optimal rational approximation [16] [^13] For example, the pole distribution pattern in the (near-)optimal rational approximation can be understood as the result of minimizing the approximation errors [16].. Fig. 3 shows such relative errors for the emulated $s$-wave $\mathcal{A}(E)$s in the complex $E$ plane. Here, the two-body binding energy $B_{2}$ is $10$ MeV and $E_{\mathrm{rel}}=10$ MeV in the sources. In the left sub-figure, the training points are lined on one black line above the real axis, but in the right, they are distributed on two black lines corresponding to a complex conjugate pair. The latter setup was mentioned previously at the end of Sec. III.2. Four emulations are contrasted within each sub-figure, based on different combinations of $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ and $N_{b}$ values.
@@
The general theme of the left sub-figure is that the errors are in the same order as those of the training calculations in the interpolation region (colored dark blue) between the training points or close to them. The errors increase when evaluating $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ (i.e., the emulated $\mathcal{A}$) further away. The orange dots representing the largest emulation errors are the locations of the DC poles of $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ because $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ diverges at those locations but the exact results are finite. Those orange dots visually trace out a BCT of the exact $\mathcal{A}$ that is significantly below the real $E$ axis, even though the high-fidelity calculation with the Hermitian $H$ operator would have a BCT along the real axis (see, e.g., Fig. 1). I.e., those DC poles form a proxy for the BCT of the exact $\mathcal{A}$ that the emulator approximates. This is the same as how the (near-)optimal rational approximation works. I now consider $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ effectively having the discretized BCT.
As it turns out, choosing an exact $T$-matrix to be compared with the emulated one could be subtle in the 4th quadrant of the complex $E$ plane. Note in the current discussion, I define the Riemann sheet of the exact $\mathcal{A}$ according to the BCT defined on the real axis (e.g., as in the analytic formulas in Appendix B). The exact result in the quadrant 1 to 3 is the one on the 1st sheet. In quadrant 4, above the new discretized BCT, the exact result used for computing emulation errors is the $\mathcal{A}$ defined on the 2nd sheet; however, below the new BCT, the exact result is the one on the 1st sheet. The ambiguity could arise for the Riemann sheet assignment in the region close to the BCT poles. Here, I connect the adjacent points by a straight line to form a continuous BCT curve, which determines the Riemann sheet of the exact $\mathcal{A}$. The well-behaved error plot near the new BCT shows the simple procedure is sufficient in the current cases; otherwise, sharp edges could show up in the region away from the discretized BCT.
@@
It is worth noting that when approaching the branch point, the emulation error increases quickly. However, the problematic region around the branching point shrinks if the approximation improves—here by reducing $\mathrm{Im}\,E_{\alpha}$ and increasing $N_{b}$. This is a general behavior of rational approximations of $\mathcal{A}$, as mentioned in Sec. II.3.
@@
In the right sub-figure, where the emulators are trained on both $E^{\mathrm{tr}}_{\alpha}$ and their complex conjugates, the dark orange dots, i.e., the discretized BCT, are back on the real axis in all the panels. This reinforces the relation between the error pattern and the BCT location. The error pattern shares the same mirror symmetry with respect to the real axis as the locations of the training points. Another possibility for the discretized BCT, while respecting the same symmetry, is to have poles distributed symmetrically on both sides of the real axis. As argued in the Appendix A, this case would have more significant emulation errors in general than the existing one with a single discretized BCT on the real axis.
@@
The difficulty in describing continuum physics with BCT on the real axis is also exposed clearly in the right sub-figure. Reducing $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ and increasing $N_{b}$ doesn’t improve $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ on the real axis, although it does improve in the region of the bound states. Therefore, one should follow the training point setup used in the right sub-figure to infer the discrete excited states from CEE; however, in the following parts of the paper focusing on continuum physics, I distribute the training points as in the left sub-figure.
@@
Figure 4: The left panels show the compressed spectra. The right ones plot the residues of the poles (see Eq. (36)) vs. the absolute values of the corresponding eigenvalues. Their insets show the absolute values of the eigenenergies in the $\log$ scale. The physical (bound) state is marked with a black “++” in the left panels and the insets of the right panels. From the top and bottom rows, $N_{b}$ increases from $3$ to $23$. Each row compares two emulation results with different $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$.
@@
Figure 4 shows the information about the compressed spectra and the residues of the corresponding poles in $\begin{bmatrix}\mathcal{A}(E)\end{bmatrix}$s (see Eq. (36)) from the emulators shown in the left sub-figure of Fig. 3. From top to bottom, $N_{b}$ increases from 3 to 23. In each panel, the results from two emulators with different $\mathrm{Im}\,(E_{\alpha}^{\mathrm{tr}})$ are compared.
@@
The left panels again show the BCT pattern of the DC state eigenenergies, while the bound state (not visible in Fig. 3) is separated from the DC states. This feature is similar to the eigenenergy distributions in the complex scaling calculations [13], but the number of eigenvalues here is much smaller than that in a typical complex scaling calculation. Moreover, in the complex scaling method, the angle between the rotated BCTs and the real axis is a control parameter, the maximum of which could be limited due to numeric difficulties. In the CEE, the moved BCTs, which emerge from spectrum compression, could be affected by the $\mathrm{Im}\,E_{\alpha}^{\mathrm{tr}}$: the smaller the $\mathrm{Im}\,E_{\alpha}^{\mathrm{tr}}$ gets, the further away the BCTs are pushed from the real axis. Such behavior is reflected in the left panels.
@@
This could be understood intuitively: reducing $\mathrm{Im}\,E_{\alpha}^{\mathrm{tr}}$ decreases the emulation errors below the real axis, and thus effectively, the failure region, where the discretized BCTs are located, is pushed further away from the real axis. Note the physical states, such as the bound state in the left panels, have less sensitivity numerically—supposedly no sensitivity at all in exact calculations.
In fact, such sensitivity difference could be exploited to separate the physical states from the DC states, in addition to relying on the distribution pattern of the eigenenergies or poles. Both ways of separating states all trace back to the fact that the discretized BCT poles represent the BCTs of the exact $\mathcal{A}$ and thus can be “redefined” without changing the function behavior above both old and new BCTs. This is also why one can use $\mathrm{Im}\,\mathcal{A}$’s behavior on the real axis, specifically its decomposition into the smooth and peak components, to “measure” resonance properties (absent here but present in the $p$-wave case). This point is further elaborated in Sec. V.
@@
The right panels in Fig. 4 plot the sizes of the pole residues ($W_{\beta}$ in Eq. (36)) against the absolute values of the eigenvalues. One can see that incredibly close to the branch point (threshold), there could be poles with tiny residues when $N_{b}$ becomes too large, such as in the $N_{b}=23$ case with over-sampled trainings. They could be the so-called Froissart doublets (or spurious poles, see Sec. 8.1 in Ref. [16]), whose contributions are negligible unless $E$ gets exceptionally close to the threshold (much smaller than $10^{-10}$ MeV here). If the thresholds are inferred from subsystem mass computations, one can separate and eliminate these spurious poles in Eq. (36). Or, as mentioned in Sec. III.4, I could apply the so-called greedy algorithm [107, 20] to eliminate redundancy, offering an effective regularization. For now, $N_{b}=3$ is under-spanning the subspace, while $N_{b}=23$ is over. Therefore, the effective dimension of the subspaces for both emulators is between $3$ and $23$.
@@
Each inset plots the absolute values of the eigenenergies on a logarithmic scale, with the bound state marked in black ++ (the number of symbols gives $N_{\mathrm{eff}}$). The distribution of the DC states in terms of the distance between poles and the branch point is quasi-even on the log scale, with a gradual tapering (i.e., density reduction) towards both the threshold and the infinity, which is already visible in the left panels with linear scales. This distribution is similar to the tapered exponential clustering of poles in the (near-)optimal rational approximations discussed in Sec. II.3. I thus identify the first similarity suggesting a close connection between CEE’s dimension reduction efficiency and the optimality of the rational approximation.
@@
Also note that when reducing $\mathrm{Im}\,E_{\alpha}^{\mathrm{tr}}$, the smallest distance between the BCT poles (or DC state eigenvalues)—excluding the spurious poles—and the threshold decreases; the distribution of poles or the eigenvalues also gets denser. The scale of that smallest separation controls to what proximity towards the threshold the emulator can be trusted because those close poles are responsible for the near-threshold behavior of $\begin{bmatrix}\mathcal{A}\end{bmatrix}$. The emulations and rational approximations eventually fail if $E$ gets too close to the threshold, as noted in the discussion of Fig. 3 and in Sec. II.3.
@@
Of course, the emulation performance is also related to the number of effective poles (excluding the spurious poles), i.e., $N_{\mathrm{eff}}$. As one can imagine, when $N_{\mathrm{eff}}$ increases and approaches the dimension of the solution subspace, the quality of near-threshold approximation saturates, and the smallest distance between the pole and branch point becomes “stuck”. (This saturation behavior is further illustrated later in the Fig. 5.) Therefore, by counting the number of symbols in the insets, i.e., the effective poles, in the $N_{b}=23$ cases, one sees that the $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}=3$ subspace is higher dimensional than that of $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}=30$. This indicates $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ could be viewed as a continuum resolution indicator for the solution manifold mapped out by varying the $\mathrm{Re}\,E$ parameter. Such resolution, an intrinsic property of the solution manifold, can only be “measured” by the emulator when its $N_{\mathrm{eff}}$ saturates the subspace dimension.
@@
Figure 5: The emulation errors vs $\sqrt{N_{\mathrm{eff}}}$. The left panels show the emulation errors for the bound state pole location and its residues. The right ones show the absolute errors for $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ (1) on the real energy axis with different $E$ values marked in the legends, and (2) averaged over the interpolation region around the training energies (marked as “int.”). From the top to bottom row, $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ reduces from $30$ to 33 MeV.
@@
Fig. 5 shows detailed information about the emulation errors vs $\sqrt{N_{\mathrm{eff}}}$ for the pole properties, including its locations and residues [^14] These residues could be related to the asymptotic normalization coefficients of the bound states [112]. in the left panels. The right panels show the same information for $\begin{bmatrix}\mathcal{A}(E)\end{bmatrix}$ in the interpolation region (marked as “int” in the legend) and along the real energy axis with locations marked in the legend (considered as extrapolations). $\mathrm{Im}\,E_{\alpha}^{\mathrm{tr}}$ decreases from the top to bottom panels. To get the data points for these plots, I run emulations with different $N_{b}$ and compute the corresponding $N_{\mathrm{eff}}$ as already described.
@@
One can see the $e^{-c\sqrt{N_{\mathrm{eff}}}}$ convergence behavior for all the quantities of my interests. Towards the largest $N_{\mathrm{eff}}$, the employed SVD-based regularization could also contribute to the emulation errors. Moreover, the extrapolation errors are generally more significant than the interpolation errors but still satisfy the same scaling law. It is also clear that reducing $\mathrm{Im}\,E_{\alpha}$ increases the maximum $N_{\mathrm{eff}}$ value that can be achieved, which is essential for lowering both interpolation and extrapolation errors. This supports the reasoning about $N_{\mathrm{eff}}$ saturating the subspace dimension during the discussion of Fig. 4. The convergence behavior is the second similarity that suggests the connection between the CEE and (near-)optimal rational approximations.
@@
Now, I look into the $p$-wave channel to investigate emulation performance with the presence of a resonance. Here, $\Lambda_{2}=200$ MeV and $\lambda=0.8$ with a resonance located at E=3.92.6iE=3.9-2.6\,i MeV. Figs. 6 and 7 show the emulation error and spectrum in the complex-$E$ plane, in parallel to the Figs. 3 and 4 for the $s$-wave. Since the resonance pole is well reproduced by the emulator, no enhanced error (i.e., dark orange dot) can be found at the resonance location in Fig. 6. The other results concerning emulation errors vs $\sqrt{N_{\mathrm{eff}}}$ are also qualitatively similar to those in the $s$-wave case, which can be found in the SM [110]. The key difference, though, is that the physical pole is now located below the real axis and thus has bigger errors than in the previous $s$-wave results. However, all the other observations are the same as in the $s$-wave case.
@@
Figure 6: Similar to Fig. 3, but for the $p$-wave case with $\Lambda_{2}=200$ MeV and $\lambda=0.8$. The resonance pole is located at 3.92.6i3.9-2.6\,i MeV.
@@
Figure 7: Similar to Fig. 4, but for the $p$-wave case with $\Lambda_{2}=200$ MeV and $\lambda=0.8$. The resonance pole is located at 3.92.6i3.9-2.6\,i MeV.
IV.3 Emulation in ($\mathrm{Re}\,E$, $\mathrm{Im}\,E$, $E_{\mathrm{rel}}$, $\lambda$, $\Lambda_{2}$)